OpenStudy (anonymous):

Write the polynomial 4x^2 + x^5 - 12 + 6x in standard form. A. 4x^2 + x^5 - 12 + 6x B. x^5 + 4x^2 +6x - 12 C. -12 +6x 4x^2 + x^5 D. x^2(4 + x^3) + 6(x - 2)

OpenStudy (anonymous):

\[x^5 + 4x^2 + 6x -12\] Standard form means that the terms are in order from highest to lowest degree.

OpenStudy (anonymous):

so Answer is B :D !

OpenStudy (anonymous):

seems pretty self-explanatory. To put an equation into standard form (depending on your definition of standard form), all you have to do is arrange the terms (meaning anything that is separated by a plus or minus) in terms of highest degree. In this case we have the terms: \[{4x^2}\] \[x^5\] \[-12\] and \[6x\] our highest degree is the term with the highest power. You can see in the above terms that \[x^5\] has the highest power; therefore, it will be first. Can you see where the rest of antoni's terms came from? x^5 is power 5 4x^2 is power 2 6x is power 1 and -12 is power 0. If you don't understand what I mean by power 0, you may not have learned that concept yet. You will learn that x^0=1 so -12*x^0=-12*1=-12. Regardless, your answer will go from power 5 to 2 to 1 to 0: standard form: \[f(x)=x^5+4x^2+6x-12\] B is your answer